OpenStudy (anonymous):

Write the expression as either the sine, cosine, or tangent of a single angle. sin 52° cos 13° - cos 52° sin 13°

jimthompson5910 (jim_thompson5910):

hint: sin(x - y) = sin(x)cos(y) - cos(x)sin(y)

jimthompson5910 (jim_thompson5910):

more trig identities like this can be found here http://www.sosmath.com/trig/Trig5/trig5/trig5.html

OpenStudy (anonymous):

@jim_thompson5910 Does it become sin65

jimthompson5910 (jim_thompson5910):

no it does not

OpenStudy (anonymous):

@jim_thompson5910 so what do it become because I really do not understand

jimthompson5910 (jim_thompson5910):

you subtract instead of add 52-13 = 39

jimthompson5910 (jim_thompson5910):

so it becomes sin(39)

OpenStudy (anonymous):

@jim_thompson5910 Thank you

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

@jim_thompson5910 can you help me with this one cos((π)/(3))cos((π)/(5)+sin((π)/(3))sin((π)/(5))

jimthompson5910 (jim_thompson5910):

now you're using the identity cos(x - y) = cos(x)cos(y) + sin(x)sin(y)

OpenStudy (anonymous):

@jim_thompson5910 π/3 - π/5?

jimthompson5910 (jim_thompson5910):

good, combine and simplify

OpenStudy (anonymous):

@jim_thompson5910 cos(2π)/(15)

jimthompson5910 (jim_thompson5910):

if you mean \[\Large \cos\left(\frac{2\pi}{15}\right)\] then you are correct

OpenStudy (anonymous):

@jim_thompson5910 thank you again

jimthompson5910 (jim_thompson5910):

yw